OpenStudy (abdela25):

HELP HELP HELP!!!! PLEASE!!! MEDALS WILL BE REWARDED TO FIRST AND BEST ANSWER !!!

OpenStudy (abdela25):

WRITE an equation in slop-intercept form for the line that passes through the given point (2,3), (1,5)

OpenStudy (anonymous):

The slope of the line would be \[\frac{ 3-5 }{2-1 } = -2\] Now, \[y-3=-2(x-2)\]\[y = -2x + 4 +3\]\[y=-2x + 7\]